LECTURE SCHEDULE, MAT 125, Spring 2005

Week of  Sections  Holidays and Exams

1/24  1.1, 1.2, 1.3
1/31  1.5, 1.6, 2.1
2/7  2.2, 2.3
2/14  2.4, review
2/21 2.5, 2.6 First Exam, Mon Feb 21, 8:30pm
2/28  2.7, 2.8, 2.9
3/7  2.10, 3.1
3/14  3.2, 3.4
3/21  Spring Break, no classes
3/28 review, 3.5 Second Exam, Tue Mar 29, 8:30pm
4/4  3.6, 3.7
4/11  4.1, 4.2
4/18  4.3, 4.5
4/25  4.6, 4.7
5/2 4.8, review Mon May 9 is last day of classes

5/9

5/16  Final, Mon May 16, 11-1:30

Rooms for the midterms and finals will be announced in class.

HOMEWORK, MAT 125, Spring 2005

Section: Topic  Problems:
1.1  Four ways to represent a function  2,18,19,36,38
1.2  Mathematical Models  4,9,11,
1.3  New functions from old  18,31,35,40,44,50
1.5  Exponential functions  10,15,17,23
1.6  Inverse functions  10,24,32,40,57
2.1  Tangent and velocity problems  4,7
2.2  Limit of a function  4,6,10,11,18
2.3  Calculating limits by limit laws  4,7,14,19,20,31
2.4  Continuity  4,14,27,31,32
2.5  Limits involving infinity  4,14,19,23,28,35,36,43
2.6  Tangents, velocities and rates of change 3,7,13,20
2.7  Derivatives  2,6,7,10a,16,21
2.8  Derivative as a function  3,14,22,25,36,47
2.9  Linear approximations  2,7,13,14
2.10  What does f′ say about f?  6,10,12,13,18,24,25
3.1  Polynomials and exponential functions 6,11,18,38,43,46,50
3.2  Product and quotient rules  6,11,16,20,28,31,36
3.4  Trigonometric functions  4,8,12,17,28,29
3.5  Chain rule  9,19,25,49,54,56,68
3.6  Implicit differentiation  5,7,10,17,23,30,54
3.7  Logarithmic functions  7,14,20,29,38
4.1  Related rates  7,9,12,14,17
4.2  Maximum and minimum values  6,24,29,40
4.3  Shapes of curves  6,7,17,28,48
4.5  l’Hospital’s rule  5,11,15,21,27,51
4.6  Optimization  2,5,10,11,13,14
4.8  Newton’s method  4,6,8,13,23